Consider the following recent financials for XYZ Corporation: Income Statement Balance Sheet Sales 73,802 Assets 209,087 Debt 38,278 Costs 44,281 Equity 170,809 EBIT 29,521 Taxes ( 38% 11,218 Total 209,087 Total 209,087 Net Income 18,303

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2,907 was paid, and the company wishes to maintain a constant payout ratio. Next year’s sales are projected to grow by 25%.

1. What is the pro-forma value for equity? (Round answer to 2 decimal places. Do not round intermediate calculations. Also, do not calculate the numbers given in the income statement and balance sheet, such as the Taxes and Net income. Take them as given.).

2. What is the external financing needed using the pro-forma approach? (Round answer to 2 decimal places. Do not round intermediate calculations. Also, do not calculate the numbers given in the income statement and balance sheet, such as the Taxes and Net income. Take them as given.).

3. What is the internal growth rate? (Report answer in percentage terms and round to 2 decimal places. Do not round intermediate calculations).

4. What is the sustainable growth rate? (Report answer in percentage terms and round to 2 decimal places. Do not round intermediate calculations).
Answer #1

Answer 1:

Pro-forma value for equity = Equity Previous year + Profit margin ratio * Sales previous year * (1 + 25%) * (1 - dividend payout ratio)

= 170809 + (18303 / 73802) * 73802 * (1 + 25%) * (1 - 2907 /18303)

= $190,054.00

Pro-forma value for equity = $190,054.00

Answer 2:

Pro-forma asset = (Asset previous year / Sales previous year) * (Sales previous year * (1 + 25%))

= (209087 / 170809) * (170809 *(1 + 25%)

= $261,358.75

External financing needed = Pro-forma asset + Debt - Pro-forma equity = 261358.75 - 38278 - 190054 = $33,026.75

Calculations are also given below:

Pro-forma Balance sheet $38,278.00 $190,054.00 Total assets $261,358.75 Total liabilities and equity $228,332.00 Assets $261,358.75 Debt quiiy External financing needed $33,026.75

External financing needed = $33,026.75

Answer 3:

Internal growth rate = RoA * Retention ratio

= Net Income / Assets * (1 - dividend payout ratio)

= 18303 /209087 * (1 - 2907 /18303)

= 7.36%

Internal growth rate = 7.36%

Answer 4:

Sustainable growth rate = RoE * Retention ratio

= Net Income / Equity * (1 - dividend payout ratio)

= 18303 /170809 * (1 - 2907 /18303)

= 9.01%

Sustainable growth rate = 9.01%